OpenStudy (anonymous):

Help?? Determine how many slices you will cut your pizza into. You need to pick two different numbers of slices. Once you have determined the number of slices, calculate the following for both ways of cutting your pizza: the interior angle and arc length of the slices the area of each slice

OpenStudy (anonymous):

Determine the delivery radius for your shop. Draw a point on a coordinate plane where your shop will be located. Create two different radii lengths from your shop, and construct the circles that represent each delivery area. How much area will each delivery radius cover? Write the equation for each circle created.

OpenStudy (anonymous):

I want one slice to be 8 and another 6 I guess? I don't know what to do afterwards.

OpenStudy (anonymous):

@modphysnoob do you know how to do it? :/ I honestly don't know...

OpenStudy (anonymous):

lol, there is a lot of read here

OpenStudy (anonymous):

true...

OpenStudy (anonymous):

can you post your full question, I am trouble making sense of this

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

First things first! You decide to focus on one size of pizza when opening your restaurant. Once you establish a good customer base, you can expand and offer additional sizes. Decide whether you’re going to focus on creating a small, medium, large, or extra-large pizza and what the diameter of that size pizza should be. Conduct some research about pizza sizes using the Internet or by calling stores in your area. Record the diameters you found during your research, and make yours slightly larger to increase competition!

OpenStudy (anonymous):

and after that question, it's the first post I posted then the second post. I'm supposed to make two slices of pizza and I guess two different numbered slices.

OpenStudy (anonymous):

ok , I get , I will do 6 slices one for you , and you do 8 slices

OpenStudy (anonymous):

do I divide that from 360?

OpenStudy (anonymous):

by*

OpenStudy (anonymous):

so I will assume pizza is 18 inch diameter ok

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

circumenference = Pi D= PI 18 now you are slicing it into 6 pieces, so arc lenght would be , pi 18/6 = pi *3

OpenStudy (anonymous):

9.42?

OpenStudy (anonymous):

3 = 9.42? :/

OpenStudy (anonymous):

is that pi times 18?

OpenStudy (anonymous):

\[\frac{9^2*\pi}{6}\]

OpenStudy (anonymous):

42.39?

OpenStudy (anonymous):

yeah that sound right area is same way r=18/2 so r=9 pi r^2= pi 18^2 divide by 6 pi 9^2 --------- 6

OpenStudy (anonymous):

42.39?

OpenStudy (anonymous):

yes, that's close

OpenStudy (anonymous):

now what do I do?

OpenStudy (anonymous):

interior angle is 360/6=

OpenStudy (anonymous):

60 for each slice

OpenStudy (anonymous):

yep, do that for 8 slices

OpenStudy (anonymous):

45 for 8 60 for 6

OpenStudy (anonymous):

yep, do area and arc lenght too

OpenStudy (anonymous):

ummm...

OpenStudy (anonymous):

is 45 and 60 the radius or just the arcs?

OpenStudy (anonymous):

what's the formula for circle area?

OpenStudy (anonymous):

pi r^2

OpenStudy (anonymous):

hmmm...

OpenStudy (anonymous):

I really don't know :/ I forgot how to get the radius...

OpenStudy (anonymous):

radius= diameter/2 we said diameter to be 18

OpenStudy (anonymous):

oh!!

OpenStudy (anonymous):

28.26

OpenStudy (anonymous):

so would it be 28.26/6 then 28.26/8?

OpenStudy (anonymous):

6 = 4.71 8 = 3.5325

OpenStudy (anonymous):

so the area would be 28.26??

OpenStudy (anonymous):

punch pi *(81/8) into calculator

OpenStudy (anonymous):

31.7925

OpenStudy (anonymous):

where did 81 come from??